evegoforth:

)A local store has a puppy that weighs 88 ounces. How many pounds does the puppy weigh, given that 1 pound = 16 ounces? (Input only the numeric values and decimal points, such as 6.4.)

jhonyy9:

so than you know the rule of 3 simple you can write 1 pound mean 16 ounces x pound mean 88 ounces and now cross multiplie and calcule the x value x = 88*1/16 = 88/16 = ? pounds hope helped understandably easy